# Selmer stability in families of congruent Galois representations

### v1

@@ 2.3. Modular Galois representations and level raising
Associated with $f$ is a modular Galois representation $\rho_{f}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}_{\mathcal{K}}(V_{f})\xrightarrow{\sim}\operatorname{GL}_{2}(\mathcal{K})$, where $V_{f}\simeq\mathcal{K}^{2}$ is the underlying vector space. There exists a Galois stable lattice $T_{f}\subset V_{f}$ and let $A_{f}:=V_{f}/T_{f}\simeq(\mathcal{K}/\mathcal{O})^{2}$. Denote the integral representation by $\rho_{T_{f}}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathcal{O})$. Let $\overline{\rho_{T_{f}}}:=\rho_{T_{f}}\pmod{(\varpi)}$ be the mod-$(\varpi)$ reduction of $\rho_{T_{f}}$.
@@
Condition (i) implies that the local representation $\rho_{|p}$ is crystalline. Condition (ii) ensures that $T_{f}$ is, up to scaling, the unique Galois-stable lattice in $V_{f}$. In light of this, and with no risk of confusion, I shall henceforth write $\rho_{f}$ (respectively, $\overline{\rho_{f}}$) in place of $\rho_{T_{f}}$ (respectively, $\overline{\rho_{T_{f}}}$), thereby omitting the explicit reference to the choice of lattice. Condition (iii) implies that $\operatorname{det}\rho_{f}=\chi$. Condition (iv) implies that $\rho_{|p}=\left({\begin{array}[]{cc}\chi\gamma&\ast\\ 0&\gamma^{-1}\\ \end{array}}\right)$ where $\gamma:\operatorname{G}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$ is a finite order unramified character.
@@
$\bar{\rho}:\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(\bar{\mathbb{F}}_{p})$
@@
The set of levels $N_{f}$ satisfying the conditions in Theorem 2.2 is denoted by $\mathcal{S}(\bar{\rho})$. For $X>0$, set $n(\bar{\rho};X):=\#\{N_{f}\in\mathcal{S}(\bar{\rho})\mid N_{f}\leq X\}$.
@@
$\bar{\rho}:\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(\bar{\mathbb{F}}_{p})$
be an irreducible Galois representation. Suppose $\bar{\rho}$ arises from a newform $g$ on $\Gamma_{1}(N_{\bar{\rho}})$ of weight $k$ with $22\leq k\leq p-$ 2, and let $M\in\mathcal{S}(\bar{\rho})$. Then there exists a cuspidal newform $f$ of weight $k$ and level $M$ such that $\bar{\rho}_{f}\simeq\bar{\rho}$.
@@ 2.4. Selmer groups
Associated to a Hecke eigenform $f$, there are two natural Selmer groups attached to its $p$-adic Galois representation. The first is defined using Greenberg’s local condition at $p$, together with the usual unramified conditions away from $p$. This Selmer group is defined over $\mathbb{Q}$, and will differ from its Iwasawa theoretic counterpart, which is defined over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$. I will refer to this as the $\mathbb{Q}$-Greenberg Selmer group of $f$. The second is the Bloch–Kato Selmer group, which is defined via the framework of $p$-adic Hodge theory.
@@
$T_{f}^{+}:=T_{f}\cap V_{f}^{+}\text{ and }T_{f}^{-}:=\pi(T_{f})$
@@
$0\rightarrow T_{f}^{+}\rightarrow T_{f}\rightarrow T_{f}^{-}\rightarrow 0$
Likewise, I set $A_{f}^{\pm}:=V_{f}^{\pm}/T_{f}^{\pm}$ and obtain the exact sequence
$0\rightarrow A_{f}^{+}\rightarrow A_{f}\rightarrow A_{f}^{-}\rightarrow 0$
Given a rational prime $\ell\neq p$, define $H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}):=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f})$. Define
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{p},A_{f}):=\operatorname{ker}\left\{H^{1}(\mathbb{Q}_{p},A_{f})\rightarrow H^{1}(\operatorname{I}_{p},A_{f}^{-})\right\}.$
@@
$\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q}):=\operatorname{ker}\left\{H^{1}(\mathbb{Q},A_{f})\rightarrow\bigoplus_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})}\right\}.$
I now define the Bloch–Kato local conditions. For any prime $\ell\neq p$, set
$H^{1}_{f}(\mathbb{Q}_{\ell},A_{f}):=\operatorname{Im}\left\{H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},V_{f})\longrightarrow H^{1}(\mathbb{Q}_{\ell},A_{f})\right\}.$
From the commutativity of the diagram
${H^{1}(\mathbb{Q}_{\ell},V_{f})}$${H^{1}(\operatorname{I}_{\ell},V_{f})}$${H^{1}(\mathbb{Q}_{\ell},A_{f})}$${H^{1}(\operatorname{I}_{\ell},A_{f})}$
it follows that $H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})$ is contained in $H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f})$. Moreover, by [Rub00, Lemma 1.3.5(i)], the subspace $H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})$ is the divisible part of $H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f})$. In particular, the index
$[H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f}):H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})]$
is finite.
Definition 2.6.
For $\ell\neq p$, the Tamagawa number of $A_{f}$ at $\ell$ is defined by
$c_{\ell}(A_{f}):=[H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f}):H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})].$
Let $\mathbf{B}_{\operatorname{cris}}$ be Fontaine’s crystalline period ring. I define the local condition at $p$ as follows:
$\begin{split}H^{1}_{f}(\mathbb{Q}_{p},V_{f})&:=\ker\left\{H^{1}(\mathbb{Q}_{p},V_{f})\longrightarrow H^{1}(\mathbb{Q}_{p},V_{f}\otimes_{\mathbb{Q}_{p}}\mathbf{B}_{\operatorname{cris}})\right\},\\ H^{1}_{f}(\mathbb{Q}_{p},A_{f})&:=\operatorname{Im}\left\{H^{1}_{f}(\mathbb{Q}_{p},V_{f})\longrightarrow H^{1}(\mathbb{Q}_{p},A_{f})\right\}.\end{split}$
Definition 2.7.
The Bloch–Kato Selmer group of $A_{f}$ over $\mathbb{Q}$ is defined by
$\operatorname{Sel}_{\operatorname{BK}}(A_{f}/\mathbb{Q}):=\ker\left\{H^{1}(\mathbb{Q},A_{f})\longrightarrow\prod_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})}{H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})}\right\}.$
The Bloch–Kato Selmer group $\operatorname{Sel}_{\operatorname{BK}}(A/\mathbb{Q})$ is contained in the Greenberg Selmer group $\operatorname{Sel}_{\operatorname{Gr}}(A/\mathbb{Q})$, see for instance [Och00, Proposition 4.1, part 1]. Further, $\operatorname{Sel}_{\operatorname{Gr}}(A/\mathbb{Q})$ is a finite index subgroup of $\operatorname{Sel}_{\operatorname{BK}}(A/\mathbb{Q})$, see [LV21, (3.1) on p.447] and [Fla90, Theorem 3].
@@ 2.5. The residual Selmer group
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi]):=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f}[\varpi]),$
@@
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{p},A_{f}[\varpi]):=\ker\left\{H^{1}(\mathbb{Q}_{p},A_{f}[\varpi])\longrightarrow H^{1}(\operatorname{I}_{p},A_{f}^{-}[\varpi])\right\}.$
@@
(2.4) 00{0}${\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})}$${H^{1}(\mathbb{Q},A_{f}[\varpi])}$${\bigoplus_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f}[\varpi])}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])}}$00{0}${\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]}$${H^{1}(\mathbb{Q},A_{f})[\varpi]}$${\bigoplus_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})[\varpi]}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})[\varpi]}.}$$\scriptstyle{\alpha}$$\scriptstyle{\Phi}$$\scriptstyle{g}$$\scriptstyle{h}$
@@
$h_{\ell}:\frac{H^{1}(\mathbb{Q}_{\ell},A_{f}[\varpi])}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])}\longrightarrow\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})[\varpi]}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})[\varpi]}.$
I set $\beta_{\ell}(A):=\dim_{\kappa}\operatorname{ker}h_{\ell}$.
@@
Let me begin by proving part (1). Consider the commutative diagram
@@
The vertical maps are injective, and therefore, $\operatorname{ker}h_{\ell}\subseteq\operatorname{ker}g_{\ell}$. From the Kummer sequence,
$\operatorname{ker}g_{\ell}=\left(\frac{H^{0}(\operatorname{I}_{\ell},A_{f})}{\varpi H^{0}(\operatorname{I}_{\ell},A_{f})}\right)^{\operatorname{G}_{\ell}/\operatorname{I}_{\ell}},$
@@
Next, I prove (2). Since the action of $\operatorname{I}_{\ell}$ on $A_{f}$ is trivial, I find that $H^{0}(\operatorname{I}_{\ell},A_{f})=A_{f}$. Since $A_{f}$ is $\varpi$-divisible, $A_{f}/\varpi A_{f}=0$ and thus part (2) follows from (1).
@@
is injective. By the Kummer sequence, its kernel is isomorphic to $\frac{H^{0}(\operatorname{I}_{p},A_{f}^{-})}{\varpi H^{0}(\operatorname{I}_{p},A_{f}^{-})}$. Since the action of $\operatorname{I}_{p}$ on $A_{f}^{-}$ is trivial, and $A_{f}^{-}$ is $\varpi$-divisible, I deduce that
$\frac{H^{0}(\operatorname{I}_{p},A_{f}^{-})}{\varpi H^{0}(\operatorname{I}_{p},A_{f}^{-})}=A_{f}^{-}/\varpi A_{f}^{-}=0.$
Consequently, $h_{p}$ is injective and thus $\beta_{p}(A_{f})=0$, proving (3).∎
@@
$\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]\leq\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})+\sum_{\ell}\beta_{\ell}(A).$
@@
A similar strategy yields lower bounds for the dimension of $\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]$ over $\kappa$. Let $S$ denote the set of primes dividing $N_{f}p$. For each $\ell\in S$, define the local condition $\mathcal{L}_{\ell}:=H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])$, so that the associated Selmer group is given by
@@
We also define the dual Selmer group by setting
$\operatorname{Sel}_{\operatorname{Gr},\perp}(A_{f}[\varpi]^{*}/\mathbb{Q}):=H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},A_{f}[\varpi]^{*}),$
@@
$\begin{split}&\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]\geq\sum_{\ell\in S}\dim_{\kappa}\operatorname{ker}h_{\ell}+\dim_{\kappa}H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{p},A_{f}[\varpi])\\ -&\dim_{\kappa}H^{0}(\mathbb{Q}_{p},A_{f}[\varpi])-\dim_{\kappa}H^{0}(\mathbb{Q}_{\infty},A_{f}[\varpi])+\dim_{\kappa}{}_{S}^{2}(A_{f}[\varpi]).\end{split}$
@@
$0\longrightarrow\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})\longrightarrow\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]\longrightarrow\ker h^{\prime}\rightarrow 0$
From the Poitou–Tate sequence (2.3), I find that
$\dim_{\kappa}\operatorname{ker}h\leq\dim_{\kappa}\operatorname{ker}h^{\prime}+\dim_{\kappa}\operatorname{Sel}_{Gr,\perp}(A_{f}[\varpi]^{*}/\mathbb{Q})-\dim_{\kappa}{}_{S}^{2}(A_{f}[\varpi]).$
@@
(2.7) $\begin{split}&\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]=\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})+\dim_{\kappa}\ker h^{\prime}\\ \geq&\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})-\dim_{\kappa}\operatorname{Sel}_{Gr,\perp}(A_{f}[\varpi]^{*}/\mathbb{Q})+\sum_{\ell\in S}\dim_{\kappa}\operatorname{ker}h_{\ell}+\dim_{\kappa}{}_{S}^{2}(A_{f}[\varpi]).\end{split}$
It follows from Wiles’ formula that
@@
Combining (2.7) and (2.8), I arrive at our result. ∎

### v2

@@ 2.3. Modular Galois representations and level raising
Let $f$ be a Hecke eigencuspform and denote the integral representation by $\rho_{T_{f}}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathcal{O})$. Let $\overline{\rho_{T_{f}}}\mathrel{\mathop{\ordinarycolon}}=\rho_{T_{f}}\pmod{(\varpi)}$ be the mod-$(\varpi)$ reduction of $\rho_{T_{f}}$.
@@
Condition (i) implies that the local representation $\rho_{|p}$ is crystalline. Condition (ii) ensures that $T_{f}$ is, up to scaling, the unique Galois-stable lattice in $V_{f}$. In light of this, and with no risk of confusion, we shall henceforth write $\rho_{f}$ (respectively, $\overline{\rho_{f}}$) in place of $\rho_{T_{f}}$ (respectively, $\overline{\rho_{T_{f}}}$), thereby omitting the explicit reference to the choice of lattice. Condition (iii) implies that $\operatorname{det}\rho_{f}=\chi$. Condition (iv) implies that $\rho_{|p}=\left({\begin{array}[]{cc}\chi\gamma&\ast\\ 0&\gamma^{-1}\\ \end{array}}\right)$ where $\gamma\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$ is a finite order unramified character.
@@
$\bar{\rho}\mathrel{\mathop{\ordinarycolon}}\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(\bar{\mathbb{F}}_{p})$
@@
The set of levels $N_{f}$ satisfying the conditions in Theorem 2.2 is denoted by $\mathcal{S}(\bar{\rho})$. For $X>0$, set $n(\bar{\rho};X)\mathrel{\mathop{\ordinarycolon}}=\#\{N_{f}\in\mathcal{S}(\bar{\rho})\mid N_{f}\leq X\}$.
@@
$\bar{\rho}\mathrel{\mathop{\ordinarycolon}}\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(\bar{\mathbb{F}}_{p})$
be an irreducible Galois representation. Suppose $\bar{\rho}$ arises from a newform $g$ on $\Gamma_{1}(N_{\bar{\rho}})$ of weight $k$ with $2\leq k\leq p-2$, and let $M\in\mathcal{S}(\bar{\rho})$. Then there exists a cuspidal newform $f$ of weight $k$ and level $M$ such that $\bar{\rho}_{f}\simeq\bar{\rho}$.
@@ 2.4. Selmer groups
Associated to a Hecke eigenform $f$, there are two natural Selmer groups attached to its $p$-adic Galois representation. The first is defined using Greenberg’s local condition at $p$, together with the usual unramified conditions away from $p$. This Selmer group is defined over $\mathbb{Q}$, and will differ from its Iwasawa theoretic counterpart, which is defined over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$. We will refer to this as the $\mathbb{Q}$-Greenberg Selmer group of $f$. The second is the Bloch–Kato Selmer group, which is defined via the framework of $p$-adic Hodge theory.
@@
$T_{f}^{+}\mathrel{\mathop{\ordinarycolon}}=T_{f}\cap V_{f}^{+}\text{ and }T_{f}^{-}\mathrel{\mathop{\ordinarycolon}}=\pi(T_{f})$
@@
$0\rightarrow T_{f}^{+}\rightarrow T_{f}\rightarrow T_{f}^{-}\rightarrow 0.$
Likewise, we set $A_{f}^{\pm}\mathrel{\mathop{\ordinarycolon}}=V_{f}^{\pm}/T_{f}^{\pm}$ and obtain the exact sequence
$0\rightarrow A_{f}^{+}\rightarrow A_{f}\rightarrow A_{f}^{-}\rightarrow 0.$
Given a rational prime $\ell\neq p$, define $H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f})$. Define
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{p},A_{f})\mathrel{\mathop{\ordinarycolon}}=\operatorname{ker}\left\{H^{1}(\mathbb{Q}_{p},A_{f})\rightarrow H^{1}(\operatorname{I}_{p},A_{f}^{-})\right\}.$
@@
$\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})\mathrel{\mathop{\ordinarycolon}}=\operatorname{ker}\left\{H^{1}(\mathbb{Q},A_{f})\rightarrow\bigoplus_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})}\right\}.$
Remark 2.6.
Another natural Selmer group to consider was introduced by Bloch and Kato [BK90]:
$\mathrm{Sel}_{\operatorname{BK}}(A_{f}/\mathbb{Q})\mathrel{\mathop{\ordinarycolon}}=\ker\!\left\{H^{1}(\mathbb{Q},A_{f})\longrightarrow\prod_{\ell}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})}{H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})}\right\},$
where the local conditions $H^{1}_{f}(\mathbb{Q}_{\ell},A_{f})$ differ from the conditions $H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})$ used to define the Greenberg Selmer group except at finitely many primes. One has a natural inclusion
$\mathrm{Sel}_{\operatorname{BK}}(A_{f}/\mathbb{Q})\subseteq\mathrm{Sel}_{\mathrm{Gr}}(A_{f}/\mathbb{Q}),$
and it is known that if $A_{f}$ is $p$-crystalline, then $\mathrm{Sel}_{\mathrm{Gr}}(A_{f}/\mathbb{Q})$ is a finite-index subgroup of $\mathrm{Sel}_{\operatorname{BK}}(A_{f}/\mathbb{Q})$ (see [OCH00, Proposition 4.1(1)], [LV21, (3.1)], and [FLA90, Theorem 3]). Consequently, when the Greenberg Selmer group vanishes, so does the Bloch–Kato Selmer group.
@@ 2.5. The residual Selmer group
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},A_{f}[\varpi]),$
@@
$H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{p},A_{f}[\varpi])\mathrel{\mathop{\ordinarycolon}}=\ker\left\{H^{1}(\mathbb{Q}_{p},A_{f}[\varpi])\longrightarrow H^{1}(\operatorname{I}_{p},A_{f}^{-}[\varpi])\right\}.$
@@
[…]
@@
$h_{\ell}\mathrel{\mathop{\ordinarycolon}}\frac{H^{1}(\mathbb{Q}_{\ell},A_{f}[\varpi])}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])}\longrightarrow\frac{H^{1}(\mathbb{Q}_{\ell},A_{f})[\varpi]}{H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f})[\varpi]}.$
We set $\beta_{\ell}(A_{f})\mathrel{\mathop{\ordinarycolon}}=\dim_{\kappa}\operatorname{ker}h_{\ell}$.
@@
Let us begin by proving part (1). Consider the commutative diagram
@@
The vertical maps are injective, and therefore, $\operatorname{ker}h_{\ell}$ injects into $\operatorname{ker}g_{\ell}$. From the Kummer sequence,
$\operatorname{ker}g_{\ell}\simeq\left(\frac{H^{0}(\operatorname{I}_{\ell},A_{f})}{\varpi H^{0}(\operatorname{I}_{\ell},A_{f})}\right)^{\operatorname{G}_{\ell}/\operatorname{I}_{\ell}},$
@@
Next, we prove (2). Since the action of $\operatorname{I}_{\ell}$ on $A_{f}$ is trivial, we find that $H^{0}(\operatorname{I}_{\ell},A_{f})=A_{f}$. Since $A_{f}$ is $\varpi$-divisible, $A_{f}/\varpi A_{f}=0$ and thus part (2) follows from (1).
@@
is injective. By the Kummer sequence, its kernel is isomorphic to $\frac{H^{0}(\operatorname{I}_{p},A_{f}^{-})}{\varpi H^{0}(\operatorname{I}_{p},A_{f}^{-})}$. Since the action of $\operatorname{I}_{p}$ on $A_{f}^{-}$ is trivial, and $A_{f}^{-}$ is $\varpi$-divisible, we deduce that $h_{p}$ is injective. Consequently $\beta_{p}(A_{f})=0$, proving (3).∎
@@
$\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]\leq\dim_{\kappa}\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})+\sum_{\ell}\beta_{\ell}(A_{f}).$
@@
A similar strategy yields lower bounds for the dimension of $\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]$ over $\kappa$. Let $S$ denote the set of primes dividing $N_{f}p$. For each $\ell\in S$, define the local condition $\mathcal{L}_{\ell}\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\operatorname{Gr}}(\mathbb{Q}_{\ell},A_{f}[\varpi])$, so that the associated Selmer group is given by
@@
One defines the dual Selmer group by setting
$\operatorname{Sel}_{\operatorname{Gr},\perp}(A_{f}[\varpi]^{*}/\mathbb{Q})\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},A_{f}[\varpi]^{*}),$
@@
[…]
@@
$0\longrightarrow\operatorname{Sel}_{\operatorname{Gr}}(A_{f}[\varpi]/\mathbb{Q})\longrightarrow\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})[\varpi]\longrightarrow\ker h^{\prime}\rightarrow 0.$
From the Poitou–Tate sequence (cf. [TAY03, p. 555, line 7]), we find that
[…]
@@
[…]
It follows from Wiles’ formula [NSW08, Theorem 8.7.9] that
@@
Combining (2.5) and (2.6), we arrive at our result. ∎